### v1

# supplementary

Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities between neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with widely disparate health outcomes. However, current statistical methods are often inadequate for appropriately defining what constitutes a spatial disparity and for constructing rankings of posterior probabilities that are robust under changes to such a definition. More specifically, non-parametric Bayesian approaches endow spatial effects with discrete probability distributions using Dirichlet processes, or generalizations thereof, and rely upon computationally intensive methods for inferring on weakly identified parameters. In this manuscript, we introduce a Bayesian linear regression framework to detect spatial health disparities. This enables us to exploit Bayesian conjugate posterior distributions in a more accessible manner and accelerate computation significantly over existing Bayesian non-parametric approaches. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of mortality rates across tracheal, bronchus, and lung cancer.
1. Introduction
We propose methods for reporting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023, and references therein) and spatial variation in health outcomes play a crucial role in detecting health disparities. While the role of spatial data analysis in epidemiological investigations has been extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and further references therein), inference pertaining to disparities comprises detecting “boundaries” on disease maps that identify significant differences among neighbors. These difference boundaries delineate regions with significantly different disease mortality and incidence rates and assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often treated as one of “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches aiming to resolve this problem include, but are not limited to, developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra and gaoSpatialDifferenceBoundary2023.
In pursuing a probabilistic learning mechanism for adjacency boundaries on a map from data available as aggregates over regions, as is typical of health administration data, we recognize that administrative boundaries between adjacent regions need not represent health outcomes measured there. In somewhat contrarian manner, the more conspicuous disease mapping models assume that neighboring regions are similar to each other. This runs counter to detecting health disparities, where we seek out differences and not similarities. Therefore, modeling spatial disparities needs to balance underlying spatial similarities based upon geographic proximity with the ability to detect differences among neighbors. li2015bayesian, gaoSpatialDifferenceBoundary2023 and aiello2023detecting, for example, propose classes of models that endow the spatial random effects with a discrete probability law using Dirichlet processes. This allows for nonzero probabilities on spatial random effects being equal to each, which are computed as posterior probabilities and used to classify spatial difference boundaries by controlling a Bayesian false discovery rate (on the lines of mullerFDRBayesianMultiple2006).
We aim to obviate some challenges seen in the aforementioned approaches. For one, the computation of the Bayesian nonparametric models devised for difference boundary analysis is cumbersome with the presence of weakly identifiable parameters that impede convergence of iterative estimation algorithms. A second issue is that, while semi-parametric approaches offer some flexibility, the hierarchical Dirichlet process obscures the spatial dependence between the random effects and makes the inference sensitive to values of hyper-parameters. The analysis presented in gaoSpatialDifferenceBoundary2023 also reveal that such models may not deliver optimal sensitivity and specificity with regard to detection of boundaries. Furthermore, ascertaining difference boundaries using such models take into account only the probability of spatial effects being different from each other and not the differences between such effects. How small differences among neighboring effects translate to higher values of the posterior probabilities of the differences remain unaccounted. And the methods do not offer a defense against the argument that small differences in the values of the spatial effects of adjacent regions need not constitute a disparity.
We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold. In this regard, we extend the approaches in lu2005bayesian and fitzpatrickEcologicalBoundaryDetection2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary and demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model. Section LABEL:sec:simulation_experiments evaluates our methods using simulations over a map of U.S. counties followed by an analysis of tracheal, bronchus and lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Testing for Significant Contrasts in Bayesian Linear Regression
2.1. Conjugate Normal-Gamma Model
Consider the Bayesian linear regression model
$y=X\beta+\eta,\;\eta\sim N_{n}(0,\sigma^{2}V_{y})\;,$ (1)
where $y$ is an $n\times 1$ vector of observations, $X$ is an $n\times p$ matrix of predictors, $\beta$ is the unknown $p\times 1$ vector of corresponding regression coefficients, and $V_{y}$ is a $n\times n$ covariance matrix that is assumed to be completely specified by the analyst. A customary conjugate prior is
$\pi(\beta,\sigma^{2})=N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{0},b_{0})\;,$ (2)
where $M_{0}$ is a $p\times p$ prior correlation matrix for $\beta$ that is fixed, $m_{0}$ is a fixed $p\times 1$ vector so that $\mathbb{E}[\beta]=M_{0}m_{0}$, $a_{0}$ and $b_{0}$ are known shape and rate parameters of the inverse-Gamma prior on $\sigma^{2}$. Familiar Bayesian calculations yield the joint posterior distribution
$\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)\;,$ (3)
where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M=\left(M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X\right)^{-1}$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We study the problem of multiple pairwise comparison within the coefficients $\beta$ when $p>2$ and thus devote our attention to studying pairwise contrasts of the form $\beta_{i}-\beta_{j}$. The fundamental question is: after observing $y$, how can we make decisions about $\beta_{i}$ being different from $\beta_{j}$ for a large number of pairs $(i,j)$?
We must clarify how we intend to ascertain if $\beta_{i}$ is different from $\beta_{j}$. We are unable to simply use $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)$ since $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)=0$ for continuous models, such as (3), regardless of the observed data. Hence, it is not possible to directly evaluate the probability of $\beta_{i}=\beta_{j}$. One possibility is to use priors with point masses on the regression coefficients, such as slab-spike priors (mitchellBayesianVariableSelection1988) and several adaptations using mixture models to meaningfully evaluate $\mathbb{P}(\beta_{i}=0\,|\,y)$. For example, scottExplorationAspectsBayesian2006 investigate multiple comparisons of normal means to 00 by modeling each null hypothesis as independent Bernoulli trials with an unknown common failure probability. However, there is some difficulty in adopting a similar strategy for testing $\beta_{i}=\beta_{j}$ in the regression setting because posterior dependence among the elements of $\beta$ implies that it would be inappropriate to model the contrasts $\beta_{i}-\beta_{j}$ being zero as independent Bernoulli random variables. In other words, hypothesis testing of several contrasts on the regression coefficients is a multiple dependent hypothesis testing problem that prevents us from justifying a conditionally independent hierarchical model.
As it stands, the continuous conjugate prior under the standard Bayesian linear regression model is incompatible with difference probabilities that only consider exact equality. Changing the prior or the model sacrifices the analytical simplicity of the conjugate prior in the Bayesian linear regression setting, so we devote our attention to the third approach by considering a new class of difference probabilities. The goal here is to generate easily interpretable and informative statements about the parameters while also prioritizing model simplicity and computational cost.
2.2. Random Effects under an Augmented Linear System
The model in (1) can be criticized as overly simplistic if, for example, both spatial autocorrelation and unstructured random error substantially contribute to the true data generating mechanism. One model that attempts to address this is the BYM (Besag, York and Mollié) model, first introduced in besagBayesianImageRestoration1991, which includes both a spatially correlated and unstructured error component as part of a Poisson generalized linear model. However, simpsonPenalisingModelComponent2017 and rieblerIntuitiveBayesianSpatial2016 explain the error components are not identifiable and scaling issues complicate interpretation of the variance parameters in the BYM model. We follow the parametric form of the BYM model in rieblerIntuitiveBayesianSpatial2016 that ensures identifiable error components and imparts better interpretation. We write this BYM2 model as
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (4)
where $\phi$ is a $n\times 1$ vector of random effects, $\rho\in(0,1)$ controls the proportion of the total error attributed to the random effects’ variance structure, and $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for the parameters in (4), then it is helpful to express the joint posterior distribution in terms of the augmented model (banerjeeModelingMassiveSpatial2020a),
$\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (5)
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
${}\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)=\text{N}_{n+p}\left(\gamma_{\star}\,|\,M_{\star}m_{\star},\sigma^{2}M_{\star}\right)\times\text{IG}\left(\sigma^{2}\,|\,a_{n},b_{n}\right)\;,$ (6)
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. An alternate representation that is interpreted in terms of least squares regression expresses $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ as the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and the scale parameter $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean of the residual sums of squares for the linear model in (5).
If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior. Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$. Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=A^{-1}(X^{\mathrm{\scriptscriptstyle T}}y+(1-\rho)m_{0})-A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}.$ (7)
A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$. If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(I-H)y$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}\;.$ (8)
In (4), inference on differences within the fixed effects $\beta$ and the spatial residuals $\gamma$ are equivalent problems if $\rho$ is known. This implies that a decision framework for the Bayesian fixed effects model specified in (1) can be extended to draw inference on both fixed and random effects in the model given by (4) after conditioning on $\rho$ and $V_{\phi}$. Thus, we turn to the original problem of computing difference probabilities in the fixed effects model before extending to spatial disparities.
2.3. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ can be analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $(1-\rho)B^{-1}=U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$. Here, $U$ does not depend on $\rho$ and $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}$ is diagonal with elements $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}$and we immediately obtain $\lim_{\rho\rightarrow 0^{+}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=0_{n\times n}$. Next, $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$, where $\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)_{ii}=d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}$ which is 0 for any $\rho\in(0,1)$ if $d_{i}=0$, and converges to 0 as $\rho\rightarrow 0^{+}$ if $d_{i}>0$. Thus, $\lim_{\rho\rightarrow 0^{+}}\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)=0_{n\times n}$. These imply that if the prior is non-informative on $\beta$, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
Turning to $\rho\rightarrow 1^{-}$, we note that if $d_{i}=0$, $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\rho$, which converges to $1$ as $\rho\rightarrow 1^{-}$. If $d_{i}>0$, then $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}=0$ and $\lim_{\rho\rightarrow 1^{-}}d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}=1$. Thus, $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$, and $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=UD^{\star}U^{-1}e$, where $e=(I-H)y$. Second order moments have limits given by $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}V$, and $\lim_{\rho\rightarrow 1^{-}}\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Furthermore, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
3. A New Framework for Detecting Spatial Disparities
3.1. Connection to Classical Hypothesis Testing
Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test a point null hypotheses $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$ using the classical linear model, $Y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is a fixed but unknown $p\times 1$ vector (see, e.g., seberHypothesisTesting2003). For each $c_{k}$, this goal is accomplished by computing the test statistic $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $\widehat{\sigma}^{2}=\frac{1}{n-p}\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. The classical approach computes a p-value $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$ where $T\sim t_{n-p}$ and accounts for multiple testing by computing adjusted p-values. If the adjusted p-value is below a pre-specified significance threshold, then $H^{(k)}_{0}$ is rejected.
In the Bayesian framework, which we adopt here, we maintain the continuous conjugate prior specified by (2) and recast the decision problem to consider the posterior probabilities of absolute standardized linear combinations among elements of $\beta$ being greater than a threshold $\epsilon$. For all $k=1,\ldots,K$ and any $\epsilon>0$, we define
$v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (9)
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations and represents the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from zero. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (9) using
$\widehat{v}_{k}(\epsilon)=\frac{1}{N}\sum_{t=1}^{N}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta^{(t)}|}{\sigma^{(t)}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\right)$ (10)
for any given $\epsilon$. Using a conjugate prior, we efficiently obtain exact posterior samples of $\beta\,|\,y$ using (3). Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which may, for example, represent mean outcomes for different treatment groups in an ANOVA.
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (9) have opposite interpretation: a small p-value, say $p_{k}$, suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0. The posterior distribution of $\beta$ with a flat prior is connected to the generalized least squares estimate $\widehat{\beta}$ in the Bayesian linear regression model. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities are perfectly correlated.
Theorem 1.
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$ for any given $\epsilon>0$.
Proof.
See Appendix LABEL:append:prop1_proof.
For the linear model, the Bayesian approach of ranking $\epsilon$-difference probabilities can replicate the rankings of the classical p-values by using a flat $\beta$ prior. Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
3.2. Extension to Spatial Random Effects
We next consider the model in (4), where each observation corresponds to one areal region. We treat $V_{\phi}$ as a fixed random effects variance structure, whereas $\rho$, which controls the proportion of the total error variance attributed to the structured component, can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the idea of $\epsilon$-difference probabilities in the fixed effects model, we now conduct posterior inference on random variables of the form $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$, for any given $\epsilon>0$. For $k=1\ldots,K$, we define the conditional posterior probability
$h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right),$ (11)
where $c_{k}\in\mathbb{R}^{n+p}$ is typically a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (7).
In the context of detecting spatial disparities, consider a set of $K$ pairs of neighboring regions denoted as $L=\{(i,j):i<j,i\sim j\}$ where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then we denote the corresponding $\epsilon$-difference probability as $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. The set $\left\{h_{ij}(\epsilon;\rho)\right\}_{(i,j)\in L}$ then represents the $\epsilon$-difference probabilities of all boundaries under consideration.
According to the equivalence between the model given by (4) and the augmented linear system discussed in Section 2.2, similar insight to the proof of Theorem 1 reveals that the rankings of the conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.

### v2

@@
 We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold.
-In this regard, we extend the approaches in lu2005bayesian and fitzpatrickEcologicalBoundaryDetection2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty.
+In this regard, we extend the approaches in lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty.
 More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions.
@@
 Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$.
-Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=A^{-1}(X^{\mathrm{\scriptscriptstyle T}}y+(1-\rho)m_{0})-A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(y-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(1-\rho)XA^{-1}m_{0})$.
+Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)/\rho\cdot A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y+(1-\rho)A^{-1}\left(X^{\mathrm{\scriptscriptstyle T}}B^{-1}X+I_{n}\right)m_{0}$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}})y-(1-\rho)B^{-1}XA^{-1}m_{0}$.
 Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
@@
 A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$.
-If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}y-(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}(I-H)y$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$.
+If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$l, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)\cdot(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y/\rho$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$.
 Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain

### v3

@@
 Abstract.
-Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities between neighboring regions on a map of disease mortality or incidence rates.
-Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with widely disparate health outcomes.
-However, current statistical methods are often inadequate for appropriately defining what constitutes a spatial disparity and for constructing rankings of posterior probabilities that are robust under changes to such a definition.
-More specifically, non-parametric Bayesian approaches endow spatial effects with discrete probability distributions using Dirichlet processes, or generalizations thereof, and rely upon computationally intensive methods for inferring on weakly identified parameters.
-In this manuscript, we introduce a Bayesian linear regression framework to detect spatial health disparities.
-This enables us to exploit Bayesian conjugate posterior distributions in a more accessible manner and accelerate computation significantly over existing Bayesian non-parametric approaches.
-Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of mortality rates across tracheal, bronchus, and lung cancer.
+Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates.
+Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes.
+However, there are statistical challenges to appropriately defining what constitutes a spatial disparity and to construct robust probabilistic inference for spatial disparities.
+We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities.
+We derive exploitable analytical tractability that considerably accelerates computation.
+Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
 1. Introduction
-We propose methods for reporting spatial disparities on maps.
-Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023, and references therein) and spatial variation in health outcomes play a crucial role in detecting health disparities.
-While the role of spatial data analysis in epidemiological investigations has been extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and further references therein), inference pertaining to disparities comprises detecting “boundaries” on disease maps that identify significant differences among neighbors.
-These difference boundaries delineate regions with significantly different disease mortality and incidence rates and assists in improved geographic allocation of health care resources.
-In geographic analysis, the problem is often treated as one of “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications).
-Probabilistic model-based approaches aiming to resolve this problem include, but are not limited to, developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra and gaoSpatialDifferenceBoundary2023.
-In pursuing a probabilistic learning mechanism for adjacency boundaries on a map from data available as aggregates over regions, as is typical of health administration data, we recognize that administrative boundaries between adjacent regions need not represent health outcomes measured there.
-In somewhat contrarian manner, the more conspicuous disease mapping models assume that neighboring regions are similar to each other.
-This runs counter to detecting health disparities, where we seek out differences and not similarities.
-Therefore, modeling spatial disparities needs to balance underlying spatial similarities based upon geographic proximity with the ability to detect differences among neighbors.
-li2015bayesian, gaoSpatialDifferenceBoundary2023 and aiello2023detecting, for example, propose classes of models that endow the spatial random effects with a discrete probability law using Dirichlet processes.
-This allows for nonzero probabilities on spatial random effects being equal to each, which are computed as posterior probabilities and used to classify spatial difference boundaries by controlling a Bayesian false discovery rate (on the lines of mullerFDRBayesianMultiple2006).
-We aim to obviate some challenges seen in the aforementioned approaches.
-For one, the computation of the Bayesian nonparametric models devised for difference boundary analysis is cumbersome with the presence of weakly identifiable parameters that impede convergence of iterative estimation algorithms.
-A second issue is that, while semi-parametric approaches offer some flexibility, the hierarchical Dirichlet process obscures the spatial dependence between the random effects and makes the inference sensitive to values of hyper-parameters.
-The analysis presented in gaoSpatialDifferenceBoundary2023 also reveal that such models may not deliver optimal sensitivity and specificity with regard to detection of boundaries.
-Furthermore, ascertaining difference boundaries using such models take into account only the probability of spatial effects being different from each other and not the differences between such effects.
-How small differences among neighboring effects translate to higher values of the posterior probabilities of the differences remain unaccounted.
-And the methods do not offer a defense against the argument that small differences in the values of the spatial effects of adjacent regions need not constitute a disparity.
-We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold.
-In this regard, we extend the approaches in lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty.
+We develop statistical methods for detecting spatial disparities on maps.
+Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps.
+Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein).
+Inference on disparities proceeds from “difference boundaries” that represent significant differences among neighbors and delineate regions with significantly different disease mortality and incidence rates, which assists in improved geographic allocation of health care resources.
+In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications).
+Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
+While the aforementioned articles present an impressive range of methods, they are broadly classified into two groups.
+The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements.
+The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities.
+Both groups introduce spatial dependence and aim to balance underlying spatial similarities with the ability to detect differences among neighbors.
+Our current contribution offers a computationally scalable and analytically tractable approach to this problem.
+Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework.
 More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions.
-Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
+Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006).
+Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in ascertaining spatial disparity, and delivering improved sensitivity and specificity for difference boundary detection (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
+We exploit closed form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold.
+We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference.
 We provide an outline for the rest of the article.
 Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects.
-Section 3 develops the notion of a difference boundary and demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model.
-Section LABEL:sec:simulation_experiments evaluates our methods using simulations over a map of U.S.
-counties followed by an analysis of tracheal, bronchus and lung cancer mortality data in Section LABEL:sec:application.
+Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model.
+Section LABEL:sec:simulation_experiments evaluates our methods using simulated data over a map of U.S.
+counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application.
 Section LABEL:sec:_discussion concludes with a brief discussion.
-2. Testing for Significant Contrasts in Bayesian Linear Regression
-2.1. Conjugate Normal-Gamma Model
-Consider the Bayesian linear regression model
-$y=X\beta+\eta,\;\eta\sim N_{n}(0,\sigma^{2}V_{y})\;,$ (1)
-where $y$ is an $n\times 1$ vector of observations, $X$ is an $n\times p$ matrix of predictors, $\beta$ is the unknown $p\times 1$ vector of corresponding regression coefficients, and $V_{y}$ is a $n\times n$ covariance matrix that is assumed to be completely specified by the analyst.
-A customary conjugate prior is
-$\pi(\beta,\sigma^{2})=N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{0},b_{0})\;,$ (2)
-where $M_{0}$ is a $p\times p$ prior correlation matrix for $\beta$ that is fixed, $m_{0}$ is a fixed $p\times 1$ vector so that $\mathbb{E}[\beta]=M_{0}m_{0}$, $a_{0}$ and $b_{0}$ are known shape and rate parameters of the inverse-Gamma prior on $\sigma^{2}$.
-Familiar Bayesian calculations yield the joint posterior distribution
-$\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)\;,$ (3)
-where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M=\left(M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X\right)^{-1}$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$.
-We study the problem of multiple pairwise comparison within the coefficients $\beta$ when $p>2$ and thus devote our attention to studying pairwise contrasts of the form $\beta_{i}-\beta_{j}$.
-The fundamental question is: after observing $y$, how can we make decisions about $\beta_{i}$ being different from $\beta_{j}$ for a large number of pairs $(i,j)$?
-We must clarify how we intend to ascertain if $\beta_{i}$ is different from $\beta_{j}$.
-We are unable to simply use $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)$ since $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)=0$ for continuous models, such as (3), regardless of the observed data.
-Hence, it is not possible to directly evaluate the probability of $\beta_{i}=\beta_{j}$.
-One possibility is to use priors with point masses on the regression coefficients, such as slab-spike priors (mitchellBayesianVariableSelection1988) and several adaptations using mixture models to meaningfully evaluate $\mathbb{P}(\beta_{i}=0\,|\,y)$.
-For example, scottExplorationAspectsBayesian2006 investigate multiple comparisons of normal means to 00 by modeling each null hypothesis as independent Bernoulli trials with an unknown common failure probability.
-However, there is some difficulty in adopting a similar strategy for testing $\beta_{i}=\beta_{j}$ in the regression setting because posterior dependence among the elements of $\beta$ implies that it would be inappropriate to model the contrasts $\beta_{i}-\beta_{j}$ being zero as independent Bernoulli random variables.
-In other words, hypothesis testing of several contrasts on the regression coefficients is a multiple dependent hypothesis testing problem that prevents us from justifying a conditionally independent hierarchical model.
-As it stands, the continuous conjugate prior under the standard Bayesian linear regression model is incompatible with difference probabilities that only consider exact equality.
-Changing the prior or the model sacrifices the analytical simplicity of the conjugate prior in the Bayesian linear regression setting, so we devote our attention to the third approach by considering a new class of difference probabilities.
-The goal here is to generate easily interpretable and informative statements about the parameters while also prioritizing model simplicity and computational cost.
+2. Assessing Contrasts in Bayesian Linear Regression
 2.1. Random Effects under an Augmented Linear System
-The model in (1) can be criticized as overly simplistic if, for example, both spatial autocorrelation and unstructured random error substantially contribute to the true data generating mechanism.
-One model that attempts to address this is the BYM (Besag, York and Mollié) model, first introduced in besagBayesianImageRestoration1991, which includes both a spatially correlated and unstructured error component as part of a Poisson generalized linear model.
-However, simpsonPenalisingModelComponent2017 and rieblerIntuitiveBayesianSpatial2016 explain the error components are not identifiable and scaling issues complicate interpretation of the variance parameters in the BYM model.
-We follow the parametric form of the BYM model in rieblerIntuitiveBayesianSpatial2016 that ensures identifiable error components and imparts better interpretation.
-We write this BYM2 model as
+We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and afford better interpretation.
+We write this “BYM2” model as
 $y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
-where $\phi$ is a $n\times 1$ vector of random effects, $\rho\in(0,1)$ controls the proportion of the total error attributed to the random effects’ variance structure, and $V_{\phi}$ is a known positive definite matrix.
-For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor.
+where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix.
+For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor.
 The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
-If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for the parameters in (4), then it is helpful to express the joint posterior distribution in terms of the augmented model (banerjeeModelingMassiveSpatial2020a),
+If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
 $\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (2)
@@
 where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$.
-An alternate representation that is interpreted in terms of least squares regression expresses $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ as the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and the scale parameter $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean of the residual sums of squares for the linear model in (5).
-If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior.
-Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$.
-Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$.
-Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)/\rho\cdot A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y+(1-\rho)A^{-1}\left(X^{\mathrm{\scriptscriptstyle T}}B^{-1}X+I_{n}\right)m_{0}$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}})y-(1-\rho)B^{-1}XA^{-1}m_{0}$.
-Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
-$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}.$ (7)
-A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$.
-If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$l, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)\cdot(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y/\rho$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$.
-Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
-$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}\;.$ (8)
-In (4), inference on differences within the fixed effects $\beta$ and the spatial residuals $\gamma$ are equivalent problems if $\rho$ is known.
-This implies that a decision framework for the Bayesian fixed effects model specified in (1) can be extended to draw inference on both fixed and random effects in the model given by (4) after conditioning on $\rho$ and $V_{\phi}$.
-Thus, we turn to the original problem of computing difference probabilities in the fixed effects model before extending to spatial disparities.
+The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
+If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
+$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
+where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior on $\beta$, i.e., $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
+$\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
+These expressions are used in Section 3.2 to formulate difference boundary detection.
 2.2. Limiting cases
-The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ can be analytically derived when the prior is non-informative on $\beta$.
+The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$.
 By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$.
-Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $(1-\rho)B^{-1}=U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$.
-Here, $U$ does not depend on $\rho$ and $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}$ is diagonal with elements $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}$and we immediately obtain $\lim_{\rho\rightarrow 0^{+}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=0_{n\times n}$.
-Next, $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$, where $\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)_{ii}=d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}$ which is 0 for any $\rho\in(0,1)$ if $d_{i}=0$, and converges to 0 as $\rho\rightarrow 0^{+}$ if $d_{i}>0$.
-Thus, $\lim_{\rho\rightarrow 0^{+}}\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)=0_{n\times n}$.
-These imply that if the prior is non-informative on $\beta$, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
-Turning to $\rho\rightarrow 1^{-}$, we note that if $d_{i}=0$, $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\rho$, which converges to $1$ as $\rho\rightarrow 1^{-}$.
-If $d_{i}>0$, then $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}=0$ and $\lim_{\rho\rightarrow 1^{-}}d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}=1$.
-Thus, $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
-Further simplifications reveal that $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$, and $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=UD^{\star}U^{-1}e$, where $e=(I-H)y$.
-Second order moments have limits given by $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}V$, and $\lim_{\rho\rightarrow 1^{-}}\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$.
-Furthermore, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis).
+Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$.
+Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$.
+Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
+We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
+Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$.
+Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$.
+Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis).
 Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
+These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).
 3. A New Framework for Detecting Spatial Disparities
 3.1. Connection to Classical Hypothesis Testing
-Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test a point null hypotheses $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$ using the classical linear model, $Y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is a fixed but unknown $p\times 1$ vector (see, e.g., seberHypothesisTesting2003).
-For each $c_{k}$, this goal is accomplished by computing the test statistic $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $\widehat{\sigma}^{2}=\frac{1}{n-p}\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates.
-The classical approach computes a p-value $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$ where $T\sim t_{n-p}$ and accounts for multiple testing by computing adjusted p-values.
-If the adjusted p-value is below a pre-specified significance threshold, then $H^{(k)}_{0}$ is rejected.
-In the Bayesian framework, which we adopt here, we maintain the continuous conjugate prior specified by (2) and recast the decision problem to consider the posterior probabilities of absolute standardized linear combinations among elements of $\beta$ being greater than a threshold $\epsilon$.
-For all $k=1,\ldots,K$ and any $\epsilon>0$, we define
+We first briefly address testing for contrasts in Bayesian linear regression without spatial effects, i.e., $\gamma=0$ in (1).
+Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003).
+We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates.
+We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
+For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate posterior probabilities for the absolute standardized linear combinations among elements of $\beta$ to exceed a threshold $\epsilon$.
+Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
 $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (6)
-We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations and represents the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from zero.
+We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 00.
 If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
@@
 for any given $\epsilon$.
-Using a conjugate prior, we efficiently obtain exact posterior samples of $\beta\,|\,y$ using (3).
-Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which may, for example, represent mean outcomes for different treatment groups in an ANOVA.
-Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (9) have opposite interpretation: a small p-value, say $p_{k}$, suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0.
-The posterior distribution of $\beta$ with a flat prior is connected to the generalized least squares estimate $\widehat{\beta}$ in the Bayesian linear regression model.
-Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities are perfectly correlated.
+Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$.
+We draw exact posterior samples from the Normal-Gamma family.
+Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean outcomes for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
+Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 00, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 00.
+In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$.
+Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
 Theorem 1.
@@
 Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$.
-Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$ for any given $\epsilon>0$.
+Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
 Proof.
-See Appendix LABEL:append:prop1_proof.
-For the linear model, the Bayesian approach of ranking $\epsilon$-difference probabilities can replicate the rankings of the classical p-values by using a flat $\beta$ prior.
+See LABEL:append:prop1_proof.
 Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities.
@@
 3.2. Extension to Spatial Random Effects
-We next consider the model in (4), where each observation corresponds to one areal region.
-We treat $V_{\phi}$ as a fixed random effects variance structure, whereas $\rho$, which controls the proportion of the total error variance attributed to the structured component, can either be a fixed hyperparameter or a random variable with a prior density.
-Generalizing the idea of $\epsilon$-difference probabilities in the fixed effects model, we now conduct posterior inference on random variables of the form $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$, for any given $\epsilon>0$.
-For $k=1\ldots,K$, we define the conditional posterior probability
+We consider (1), where each observation corresponds to one region.
+We let $V_{\phi}$ be known, whereas $\rho$ can either be a fixed hyperparameter or a random variable with a prior density.
+Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we conduct posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$.
+For $k=1\ldots,K$, we define
 $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right),$ (8)
-where $c_{k}\in\mathbb{R}^{n+p}$ is typically a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
+where $c_{k}\in\mathbb{R}^{n+p}$ is a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
 If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (4).
-In the context of detecting spatial disparities, consider a set of $K$ pairs of neighboring regions denoted as $L=\{(i,j):i<j,i\sim j\}$ where $i\sim j$ means that regions $i$ and $j$ are neighbors.
-If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then we denote the corresponding $\epsilon$-difference probability as $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$.
-The set $\left\{h_{ij}(\epsilon;\rho)\right\}_{(i,j)\in L}$ then represents the $\epsilon$-difference probabilities of all boundaries under consideration.
-According to the equivalence between the model given by (4) and the augmented linear system discussed in Section 2.2, similar insight to the proof of Theorem 1 reveals that the rankings of the conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.
+In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that regions $i$ and $j$ are neighbors.
+If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$.
+Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the rankings of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.

### v4

@@
-supplementary
+Assessing Spatial Disparities: A Bayesian Linear Regression Approach
 Abstract.
 Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates.
-Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes.
-However, there are statistical challenges to appropriately defining what constitutes a spatial disparity and to construct robust probabilistic inference for spatial disparities.
+Analyzing such data introduces spatial dependence among health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes.
+However, there are statistical challenges to appropriately define what constitutes a spatial disparity and to construct robust probabilistic inferences for spatial disparities.
 We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities.
 We derive exploitable analytical tractability that considerably accelerates computation.
-Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
+Simulation experiments conducted on a county map of the entire United States demonstrate the effectiveness of our method, and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
 1. Introduction
 We develop statistical methods for detecting spatial disparities on maps.
-Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps.
+Disparities in health outcomes are a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps.
 Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein).
-Inference on disparities proceeds from “difference boundaries” that represent significant differences among neighbors and delineate regions with significantly different disease mortality and incidence rates, which assists in improved geographic allocation of health care resources.
+Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources.
 In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications).
 Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
-While the aforementioned articles present an impressive range of methods, they are broadly classified into two groups.
+The aforementioned articles present an impressive range of methods, but are broadly classified into two groups.
 The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements.
 The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities.
-Both groups introduce spatial dependence and aim to balance underlying spatial similarities with the ability to detect differences among neighbors.
+Both groups introduce spatial dependence and aim to balance the underlying spatial similarities with the ability to detect differences among neighbors.
 Our current contribution offers a computationally scalable and analytically tractable approach to this problem.
@@
 Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006).
-Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in ascertaining spatial disparity, and delivering improved sensitivity and specificity for difference boundary detection (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
-We exploit closed form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold.
+The benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in determining spatial disparity, and providing improved sensitivity and specificity for the detection of boundary differences (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
+Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in slackGeographicConcentrationUs2014; jagai2017county; rodriguez2018county).
+We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold.
 We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference.
@@
 Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model.
-Section LABEL:sec:simulation_experiments evaluates our methods using simulated data over a map of U.S.
-counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application.
+Section LABEL:sec:simulation_experiments evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application.
 Section LABEL:sec:_discussion concludes with a brief discussion.
@@
 2.1. Random Effects under an Augmented Linear System
-We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and afford better interpretation.
+We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and provide a better interpretation.
 We write this “BYM2” model as
 $y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
-where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix.
-For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor.
-The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
+where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper.
+The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
+For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis.
+Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=N(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$.
+The conditional distribution of each spatial residual is based on a linear combination of its neighbors.
+By Brook’s lemma (besagBayesianImageRestoration1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (banerjeeHierarchicalModelingAnalysis2015, Chap.
+4).
+The density $\pi(\phi)$ is proper when $\alpha\in(1/\lambda_{min},1)$, where $\lambda_{min}$ is the smallest eigenvalue of $D_{W}^{-1/2}WD_{W}^{-1/2}$.
+Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
 If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
@@
 where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$.
-The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
+The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least squares regression estimate to be solved $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
 If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
 $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
-where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior on $\beta$, i.e., $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
+where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior to $\beta$, that is, $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
 $\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
-These expressions are used in Section 3.2 to formulate difference boundary detection.
+These expressions are used in Section 3.2 to formulate the detection of differences between boundaries.
 2.2. Limiting cases
 The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$.
-By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$.
+By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$.
 Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$.
@@
 Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
-We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
-Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$.
+We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with the entries $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
+Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$ are the residuals of ordinary least squares.
 Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$.
-Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis).
+Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$.
 Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
@@
 3.1. Connection to Classical Hypothesis Testing
-We first briefly address testing for contrasts in Bayesian linear regression without spatial effects, i.e., $\gamma=0$ in (1).
-Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003).
-We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates.
+We first briefly address contrast testing in Bayesian linear regression without spatial effects, that is, $\gamma=0$ in (1).
+Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ nonzero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003).
+We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are generalized least squares estimates.
 We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
-For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate posterior probabilities for the absolute standardized linear combinations among elements of $\beta$ to exceed a threshold $\epsilon$.
+For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate the posterior probabilities for the absolute standardized linear combinations between elements of $\beta$ to exceed a threshold $\epsilon$.
 Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
 $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (6)
-We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 00.
+We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 0.
 If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
@@
 We draw exact posterior samples from the Normal-Gamma family.
-Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean outcomes for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
-Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 00, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 00.
+Specific domain applications may provide further information on possible constraints on $\epsilon$ and what is considered to be a clinically significant difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean results for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
+The classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0.
 In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$.
@@
 Theorem 1.
-Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed.
-Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$.
+Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$ and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed.
+Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value to test the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$.
 Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
 Proof.
-See LABEL:append:prop1_proof.
-Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities.
+See Appendix LABEL:append:prop1_proof.
+Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with the two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities.
 Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@
 We consider (1), where each observation corresponds to one region.
-We let $V_{\phi}$ be known, whereas $\rho$ can either be a fixed hyperparameter or a random variable with a prior density.
-Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we conduct posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$.
+We let $V_{\phi}$ be known, while $\rho$ can be either a fixed hyperparameter or a random variable with a prior density.
+Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we perform posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$.
 For $k=1\ldots,K$, we define
@@
 If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (4).
-In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that regions $i$ and $j$ are neighbors.
-If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$.
-Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the rankings of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.
+In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that the regions $i$ and $j$ are neighbors.
+If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ are such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$.
+Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the ranking of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ is stable with respect to $\epsilon$ for any $\rho\in(0,1)$.

### v5

@@
 1. Introduction
-We develop statistical methods for detecting spatial disparities on maps.
-Disparities in health outcomes are a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps.
-Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein).
-Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources.
-In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications).
-Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
+We develop statistical methods for detecting local spatial disparities on maps.
+Disparities in health outcomes are a rapidly evolving field within public health (Rao, 2023) and spatial variation in health outcomes affects how disparities manifest on disease maps.
+Spatial data analysis in epidemiological investigations is extensively documented (Waller and Gotway, 2004; Waller and Carlin, 2010; Lawson, 2013; Lawson et al., 2016).
+Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources and guide future investigation into local determinants.
+Boundary detection enables precise identification of health event clusters and sudden changes in latent or observed environmental exposures, which is crucial for defining sampling protocols across a diverse range of health environments and planning targeted interventions (Copeland, 2010; Jacquez, 2010).
+For example, Barboza-Salerno et al.
+(2025) apply boundary detection to identify local socioeconomic inequalities and estimate their association with neighborhood child mortality.
+In geographic analysis, the problem is often called “wombling” (see Jacquez and Greiling, 2003aa, 2003bb; Lu and Carlin (2005); Li et al.
+(2011); Ma and Carlin (2007); Fitzpatrick et al.
+(2010) for some algorithmic approaches with applications).
+Probabilistic model-based approaches include developments in Lu et al.
+(2007); Ma et al.
+(2010); Li et al.
+(2015, 2012); Hanson et al.
+(2015); Lee and Mitchell (2012); Corpas-Burgos and Martinez-Beneito (2020); Gao et al.
+(2023) and Aiello and Banerjee (2025).
 The aforementioned articles present an impressive range of methods, but are broadly classified into two groups.
@@
 More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions.
-Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006).
-The benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in determining spatial disparity, and providing improved sensitivity and specificity for the detection of boundary differences (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
-Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in slackGeographicConcentrationUs2014; jagai2017county; rodriguez2018county).
+Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (Müller et al., 2007).
+Controlling the false discovery rate is preferred to controlling the family wise error rate or false positive rate because it offers increased power and flexibility (Catelan and Biggeri, 2010; Glickman et al., 2014) and has been previously applied in analyses of health disparities (Tian et al., 2010; Slutske et al., 2023).
+Further developments of Bayesian FDR for multiple testing in spatial settings include Ventrucci et al.
+(2011); Sun et al.
+(2015); Tansey et al.
+(2018).
+Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects in determining spatial disparity, and providing improved sensitivity and specificity for detecting boundary differences (as seen in Gao et al., 2023).
+Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in Slack et al., 2014; Jagai et al., 2017; Rodriguez et al., 2018).
 We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold.
-We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference.
+We extend Lu and Carlin (2005) and Fitzpatrick et al.
+(2010), who use only posterior means or spatial residuals to detect differences, to fully model-based inference.
 We provide an outline for the rest of the article.
@@
 Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model.
-Section LABEL:sec:simulation_experiments evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application.
-Section LABEL:sec:_discussion concludes with a brief discussion.
+Section 4 evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section 5.
+Section 6 concludes with a brief discussion.
 2. Assessing Contrasts in Bayesian Linear Regression
 2.1. Random Effects under an Augmented Linear System
-We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and provide a better interpretation.
-We write this “BYM2” model as
-$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
+We follow the adaptation by Riebler et al. (2016) of the “BYM” model (Besag et al., 1991) to ensure identifiable parameters and provide better interpretation. The “BYM2” model is
+$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim\mbox{N}_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim\mbox{N}_{n}(0,I_{n})$ (1)
 where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper.
-The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
+The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (Riebler et al., 2016).
 For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis.
-Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=N(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$.
+Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=\mbox{N}(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$.
 The conditional distribution of each spatial residual is based on a linear combination of its neighbors.
-By Brook’s lemma (besagBayesianImageRestoration1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (banerjeeHierarchicalModelingAnalysis2015, Chap.
+By Brook’s lemma (Besag et al., 1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (Banerjee et al., 2015, Chap.
 4).
@@
 Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
-If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
+If $\pi(\gamma,\beta,\sigma^{2},\rho)=\mbox{N}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\mbox{N}(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
 $\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (2)
 where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system).
-The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
+By conjugacy of the prior $\pi(\gamma_{\star},\sigma^{2}\,|\,\rho)$ with the Gaussian likelihood in (2), the resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y_{\star},\rho)\propto\pi(y_{\star}\,|\,\gamma_{\star},\sigma^{2},\rho)\times\pi(\gamma_{\star},\sigma^{2}\,|\,\rho)$, is
 $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)=\text{N}_{n+p}\left(\gamma_{\star}\,|\,M_{\star}m_{\star},\sigma^{2}M_{\star}\right)\times\text{IG}\left(\sigma^{2}\,|\,a_{n},b_{n}\right)\;,$ (3)
-where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$.
-The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least squares regression estimate to be solved $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
-If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
+[…]
+Fixing $\rho\in(0,1)$ we can draw exact posterior samples from (3). Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
 $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
@@
 $\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
-These expressions are used in Section 3.2 to formulate the detection of differences between boundaries.
+Section 3.2 uses these expressions to formulate the detection of differences between boundaries.
 2.2. Limiting cases
@@
 By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$.
-Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$.
-Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$.
-Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
-We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with the entries $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
-Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$ are the residuals of ordinary least squares.
-Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$.
+Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, and $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$.
+The block entries of $M_{\star}$ and $\widehat{\gamma}_{\star}$ in (3) reveal that as $\rho\rightarrow 0^{+}$,
+$\begin{split}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]&=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y\longrightarrow 0_{n},\\ \operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)&=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}\longrightarrow 0_{n\times n}.\end{split}$ (6)
+Therefore, with a flat prior on $\beta$, $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)=\delta_{0_{n}}(\gamma)$(degenerate mass at $\gamma=0_{n}$).
+We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$.
+Simplifications reveal
+$\begin{split}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]&\longrightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right),\quad\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\longrightarrow UD^{\star}U^{-1}e,\\ \operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)&\longrightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1},\quad\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\longrightarrow\sigma^{2}V,\end{split}$ (7)
+as $\rho\rightarrow 1^{-}$, where e=(IH)ye=(I-H)y are the residuals from ordinary least squares and $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$.
 Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$.
 Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
-These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).
+Section 3.5 discusses their implications for boundary detection.
 3. A New Framework for Detecting Spatial Disparities
 3.1. Connection to Classical Hypothesis Testing
-We first briefly address contrast testing in Bayesian linear regression without spatial effects, that is, $\gamma=0$ in (1).
-Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ nonzero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003).
-We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are generalized least squares estimates.
-We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
-For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate the posterior probabilities for the absolute standardized linear combinations between elements of $\beta$ to exceed a threshold $\epsilon$.
+[…]
+In contrast to assuming strict equality in the point null hypothesis, equivalence tests define a margin around the null value where the effect size is not considered to be practically different from the null value.
+Bayesian approaches based on this “region of practical equivalence” (ROPE) follow a similar logic and enable full probabilistic statements about the distribution of the parameter within the ROPE (Kruschke and Liddell, 2018; Makowski et al., 2019).
+Here, we follow a similar approach by retaining the continuous conjugate prior for $\beta$ and evaluating the posterior probabilities for the absolute standardized linear combinations between elements of $\beta$ to exceed a threshold $\epsilon$.
 Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
 $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (8)
-We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 0.
+We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold or equivalence margin across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 0.
 If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (8) using
@@
 for any given $\epsilon$.
-Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$.
+Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=\mbox{N}(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$.
 We draw exact posterior samples from the Normal-Gamma family.
-Specific domain applications may provide further information on possible constraints on $\epsilon$ and what is considered to be a clinically significant difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean results for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
+Specific domain applications can provide further information on possible constraints for $\epsilon$ and on defining a clinically significant difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean results for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
 The classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (8) have opposite interpretation: a small value of $p_{k}$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0.
@@
 Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$ and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed.
-Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value to test the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$.
+Under the standard linear model $y\sim\mbox{N}_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value to test the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$.
 Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
 Proof.
-See Appendix LABEL:append:prop1_proof.
-Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with the two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities.
+See Section LABEL:append:prop1_proof.
+Theorem 1 implies that the ascending order of the p-values associated with null hypotheses $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with the two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities.
 Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@
 If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ are such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$.
+Here, $\epsilon$ represents the equivalence margin for the combined contribution from all latent health effects averaged within each area after adjusting for known risk factors in $X$.
 Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the ranking of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ is stable with respect to $\epsilon$ for any $\rho\in(0,1)$.
+Theorem 2.
+Let $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{N}_{p}(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$ be the prior distribution for the model in (1).
+Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{n+p}$ and $X\in\mathbb{R}^{n\times p}$ be fixed.
+Define $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right)$ for $k=1,\ldots,K$, where $\epsilon>0$ and $\rho\in(0,1)$.
+If there exists $\epsilon_{\star}>0$ such that $h_{k}(\epsilon_{\star};\rho)<h_{k^{\prime}}(\epsilon_{\star};\rho)$ for any $k\neq k^{\prime}$, then $h_{k}(\epsilon;\rho)<h_{k^{\prime}}(\epsilon;\rho)$.
+Proof.
+See Section LABEL:append:prop2_proof.
+If $\pi(\rho)$ denotes the prior for $\rho$, then the marginal posterior distribution of $\gamma_{\star}$ is analytically inaccessible and we define the $\epsilon$-difference probability for any $c_{k}\in\mathbb{R}^{(n+p)\times 1}$ as
+$v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}c_{k}}}>\epsilon\;\right|\;y\right)=\int_{0}^{1}h_{k}(\epsilon;\rho)\pi(\rho\,|\,y)d\rho\;.$ (11)
+Given posterior samples $\gamma_{\star}^{(1)},\ldots,\gamma_{\star}^{(N)}$ and $\rho^{(1)},\ldots,\rho^{(N)}$, we estimate (11) by
+$\widehat{v}_{k}(\epsilon)=\frac{1}{N}\sum_{t=1}^{N}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}^{(t)}|}{\sigma^{(t)}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}\left(X_{\star}^{\mathrm{\scriptscriptstyle T}}\left(V_{y_{\star}}^{(t)}\right)^{-1}X_{\star}\right)^{-1}c_{k}}}>\epsilon\right)\;,$ (12)
+where $V_{y_{\star}}^{(t)}=\begin{bmatrix}(1-\rho^{(t)})I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho^{(t)}V_{\phi}\end{bmatrix}$.
+Since $V_{y_{\star}}^{(t)}$ depends on the sampled value of $\rho^{(t)}$, computing $\operatorname{Var}(\gamma_{\star}\,|\,y,\sigma^{2},\rho)$ for each posterior sample requires inverting an $(n+p)\times(n+p)$ matrix for each posterior sample, which can become cumbersome.
+When we are only interested in linear combinations of spatial effects $\gamma$, the computation of $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is significantly accelerated by utilizing $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=U\left(I_{n}+\frac{\rho}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$ as an alternate expression for (5) in Algorithm LABEL:alg:gibbs_sampling, where $U$ and $D$ are defined in Section 2.2 with a flat prior on $\beta$.
+The choice of $\pi(\rho)$ is crucial for efficiently sampling the posterior joint of $\{\beta,\gamma,\sigma^{2},\rho\}$ using MCMC and computing (12).
+Non-informative priors, such as uniform distributions on $[0,1]$, induce extreme posterior overfitting in overly parameterized models (Gelman, 2006; Simpson et al., 2017).
+Instead, we consider the class of penalized complexity (PC) priors introduced in Simpson et al.
+(2017) as a means to encourage model parsimony and construct priors imparting a better interpretation.
+We follow Riebler et al.
+(2016) by placing a PC prior on $\rho$, denoted by $\mbox{PC}(\rho\,|\,\lambda_{\rho},V_{\phi})\propto\lambda_{\rho}\exp(-\lambda_{\rho}d(\rho;V_{\phi}))$ for $0\leq\rho\leq 1$, where $\lambda_{\rho}>0$ is a fixed hyperparameter, $d(\rho;V_{\phi})=\sqrt{2D_{KL}(p(y;\rho)||q(y))}$ and $D_{KL}(p(y;\rho)||q(y))$ is the Kullback-Leiber divergence between $p(y;\rho)=\text{N}_{n}(y\,|\,0_{n},\rho V_{\phi}+(1-\rho)I_{n})$ and the base model without structured variance $q(y)=\text{N}_{n}(y\,|\,0_{n},I_{n})$.
+The hyperparameter $\lambda_{\rho}$ is chosen such that $P(\rho\leq U)=a$ where $U$ and $a$ are constants.
+Riebler et al.
+(2016) present promising simulation results demonstrating the PC prior’s ability to shrink towards the simpler models with only spatial or non-spatial error variance.
+When $\pi(\beta,\gamma,\sigma^{2},\rho)=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\mbox{PC}(\rho\,|\,\lambda_{\rho},V_{\phi})$, Algorithm LABEL:alg:update_steps provides Gibbs updates for $\{\beta,\gamma,\sigma^{2}\}$ and a Metropolis-Hastings update for $\rho$.
+We assess predictive fit using data replicates, where one instance of $y_{rep}^{(t)}\sim\text{N}_{n}\left(X\beta^{(t)}+\gamma^{(t)},\sigma^{2(t)}(1-\rho^{(t)})I_{n}\right)$ is sampled for each drawing of the posterior samples $\{\beta^{(t)},\gamma^{(t)},\sigma^{2(t)},\rho^{(t)}\}$.
+3.3. Difference Probabilities in Generalized Linear Mixed Models
+The stability of rankings of conditional posterior probabilities is established in the previous sections using a normally distributed outcome.
+More generally, for generalized linear mixed models theoretical tractability is lost when the outcome follows a probability distribution from the exponential family with link function
+$g(\mathbb{E}(y_{i}\,|\,\beta,\gamma_{i},\eta_{i}))=x_{i}^{\mathrm{\scriptscriptstyle T}}\beta+\gamma_{i}+\eta_{i},$ (13)
+where $x_{i}$ is $p\times 1$ consisting of values of covariates for the $i$-th observation, $i=1,\ldots,n$, $\gamma\sim\mbox{N}_{n}(\gamma\,|\,0_{n},\sigma^{2}\rho V_{\phi})$, and $\eta\sim\mbox{N}_{n}(0,\sigma^{2}(1-\rho)I_{n})$.
+The link function $g(\cdot)$ relates the mean result to covariates and random effects and is commonly chosen as the canonical link such that (13) is the natural parameter of the exponential family distribution of $y_{i}$.
+We use posterior samples using MCMC to estimate difference probabilities of the form
+$\tau_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sqrt{\operatorname{Var}(c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}\,|\,y)}}>\epsilon\;\right|\;y\right),$ (14)
+where $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$.
+The marginal posterior variance is estimated with the posterior samples by $\widehat{\operatorname{Var}}(c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}\,|\,y)=\frac{1}{T-1}\sum_{t=1}^{T}(c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}-c_{k}^{\mathrm{\scriptscriptstyle T}}\bar{\gamma}_{\star})^{2}$ where $\bar{\gamma}_{\star}=\frac{1}{T}\sum_{t=1}^{T}\gamma^{(t)}_{\star}$.
+The order of $\tau_{1},\ldots,\tau_{K}$ may still be stable with respect to choice of $\epsilon$ if the marginal posterior distribution of $\gamma_{\star}$ resembles, for example, a multivariate normal or t distribution.
+We note that this approach for non-Gaussian outcomes has a weakness for detecting spatial disparities: the choice of the link function $g(x)$ in (13) critically determines the interpretation of the difference probability.
+For example, if $g(x)=\log(x)$, the canonical link for a Poisson generalized linear model, then the differences $|\phi_{i}-\phi_{j}|$ must now be interpreted in terms of unscaled difference in adjusted log-relative risk, which may produce non-intuitive results.
+The difference probabilities corresponding to the boundaries in high-risk boundaries may be ranked lower than those of low-risk regions with similar geographic heterogeneity in observed rates.
+Depending on other risk factors, disparities in log-relative risk may be more frequent in low-risk regions with minimal differences in incidence rates.
+Difference probabilities based on transformations of the spatial residuals can offer one alternative resolution but may result in a different set of reported boundaries; we demonstrate this through a simulation example in Section 4.2, where the difference probability defined in (14) achieves superior classification performance when compared to a similar difference probability defined in terms of an adjusted incidence rate ratio.
+We otherwise focus our attention on the case of a normally distributed outcome and identity link function, applying equal weight to differences between rates in high-risk and low-risk regions.
+3.4. A Bayesian FDR-Based Decision Rule.
+After computing the difference probability $v_{k}(\epsilon)$ for each linear combination of interest, we wish to account for multiple comparisons before constructing binary decisions because we are not operating under an automatically penalizing mixture model as in Scott and Berger (2006).
+We specifically safeguard against Type I errors by controlling the false discovery rate (FDR introduced in Benjamini and Hochberg, 1995), while retaining power by minimizing the false negative rate (FNR).
+Thus, we define $\text{FDR}=\frac{\sum_{k=1}^{K}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}\leq\epsilon\right)d_{k}}{\iota+\sum_{k=1}^{K}d_{k}}$, $\text{FNR}=\frac{\sum_{k=1}^{K}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)(1-d_{k})}{\iota+\sum_{k=1}^{K}(1-d_{k})}$ where $\iota>0$ is a small constant to avoid a zero denominator and $d_{k}\in\{0,1\}$ represents our decision to declare $\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon$.
+The FDR and FNR functions are random variables that have different interpretations in the two paradigms Müller et al.
+(2007).
+Given a set of difference probabilities $\{v_{1}(\epsilon),\ldots,v_{K}(\epsilon)\}$, we use the Bayesian FDR and FNR defined in Müller et al. (2004) as the posterior expectation
+$\overline{\text{FDR}}(t^{\star},\epsilon)=\mathbb{E}[\text{FDR}(t^{\star},\epsilon)\,|\,y]=\frac{\sum_{k=1}^{K}(1-v_{k}(\epsilon))\text{I}(v_{k}(\epsilon)\geq t^{\star})}{\sum_{k=1}^{K}\text{I}(v_{k}(\epsilon)\geq t^{\star})},$ (15)
+$\overline{\text{FNR}}(t^{\star},\epsilon)=\mathbb{E}[\text{FNR}(t^{\star},\epsilon)\,|\,y]=\frac{\sum_{k=1}^{K}v_{k}(\epsilon)\text{I}(v_{k}(\epsilon)<t^{\star})}{\sum_{k=1}^{K}\text{I}(v_{k}(\epsilon)<t^{\star})}.$ (16)
+Minimizing the $\overline{\text{FNR}}$ subject to $\overline{\text{FDR}}\leq\delta$ is achieved by decisions of the form $d_{k}=\text{I}(v_{k}(\epsilon)\geq t^{\star}(\epsilon))$, where $t^{\star}(\epsilon)\in[0,1]$ for all $\epsilon>0$ is a threshold chosen as
+$t^{\star}(\epsilon)=\inf\left\{\;t\in[0,1]:\overline{\text{FDR}}(t,\epsilon)\leq\delta\right\}.$ (17)
+Bayesian and classical FDR compute posterior probabilities and p-values, respectively, which are used to rank the plausibility of each hypothesis before choosing a cut-off according to a decision rule.
+In general, the Bayesian FDR and its classical analogue are fundamentally different criterion that are not necessarily equivalent, and their respective control procedures may produce different sets of decisions.
+However, in the linear fixed effects model and the setting of Theorem 1, if the prior is noninformative on $\beta$ in the Bayesian approach, then the order of the posterior probabilities is the reversed order of the p-values.
+In this case, the Bayesian and classical FDR control procedures differ only in the decision criterion; specific choice of the decision parameter $\delta$ in the Bayesian approach can afford decisions identical to many classical FDR control procedures.
+For creating binary decisions, $\epsilon$ must be chosen to apply this Bayesian FDR control procedure.
+However, for any $\delta>0$, if we set $\epsilon$ to be arbitrarily small and $t^{\star}=0$, we would declare every decision positive as $d_{k}=1$.
+On the other hand, if we set $\epsilon$ to be arbitrarily large and $t^{\star}=1$, then we would declare every decision negative as $d_{k}=0$.
+Both cases achieve $\overline{\text{FDR}}\leq\delta$, but do not induce meaningful statements about the data.
+This motivates a heuristic to choose $\epsilon$ at a reasonable value before applying the Bayesian FDR control procedure to create decisions.
+3.5. Selecting an $\epsilon$-difference threshold
+In the absence of guidance on the choice of $\epsilon$, we consider two general heuristics to select a reasonable value for $\epsilon$, the first based on a loss function and the relationship between $\epsilon$ and our resulting actions.
+As $\epsilon$ increases, the event that a contrast exceeds $\epsilon$ standard deviations gains stronger meaning, but has a lower posterior probability.
+On the other hand, the event that a contrast is less than $\epsilon$ standard deviations becomes less precise but more probable.
+This motivates the use of a loss function that is optimized with respect to $\epsilon$ by balancing the strength of our affirmative conclusions that a boundary is a disparity at the $\epsilon$-difference level with the precision of our negative conclusions.
+As a general heuristic, we appeal to the principle of maximum entropy, which suggests that, given a selection of possible posterior distributions, the posterior distribution that best represents our posterior knowledge is precisely the one with maximum entropy.
+The principle of maximum entropy (Jaynes, 1957) selects the maximum entropy distribution due to it being “maximally noncommittal” with respect to the amount of information we imbue the posterior with beyond the observed data.
+Under the model (1), choosing an extremely small or large difference threshold $\epsilon$ causes the posterior distribution of $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ to collapse into a point mass at 1 or 0 respectively, which has minimal entropy.
+Maximizing the joint posterior entropy of $R=\left\{r_{1}(\epsilon),\ldots,r_{K}(\epsilon)\right\}$ with respect to $\epsilon$ thus avoids obscuring the information that the data $y$ provides on the true quantities of interest, $\left\{\frac{|c_{1}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{1}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{1}}},\ldots,\frac{|c_{K}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{K}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{K}}}\right\}$.
+Shannon (1948) explains the Shannon entropy as an objective measure of the information or uncertainty of a probability distribution.
+Let $I$ be a $K\times 1$ vector such that $I_{k}\in\{0,1\}$ for all $k=1,\ldots,K$ and let $g(I)=\mathbb{P}(\cap_{k=1}^{K}r_{k}(\epsilon)=I_{k}\,|\,y)$.
+Since $r_{k}(\epsilon)$ is a binary random variable equal to 1 with posterior probability $v_{k}(\epsilon)$, the joint Shannon posterior entropy of $R$ measured in nats is defined as $H(R)=-\sum_{I\in\{0,1\}^{K}}g(I)\log g(I)$.
+However, the joint entropy entails estimating $2^{K}$ probabilities and is computationally unfeasible even for moderate $K$.
+Instead, we opt for maximizing the entropy of a uniformly chosen $\epsilon$-difference indicator.
+For any $\epsilon>0$, let $J$ be a random variable with uniform probability mass on $\{1,\ldots,K\}$ and $W_{\epsilon}=r_{J}(\epsilon)$.
+The conditional posterior entropy of $W_{\epsilon}$ given $J$ is $H_{J}(W_{\epsilon})=\frac{1}{K}\sum_{k=1}^{K}H(r_{k}(\epsilon))$.
+The sum of the individual entropies is related to the joint entropy by the sub-additivity property of the joint entropy (Shannon, 1948), which states that $H(R)\leq\sum_{k=1}^{K}H(r_{k}(\epsilon))$, where $H(r_{k}(\epsilon))=-v_{k}(\epsilon)\log v_{k}(\epsilon)-(1-v_{k}(\epsilon))\log(1-v_{k}(\epsilon))$.
+We obtain an optimal $\epsilon$ by maximizing the average of the individual entropies or, equivalently, minimizing
+$\text{LOSS}_{\text{CE}}(\epsilon)=\sum_{k=1}^{K}v_{k}(\epsilon)\log v_{k}(\epsilon)+(1-v_{k}(\epsilon))\log(1-v_{k}(\epsilon))\;,$ (18)
+which we refer to as the conditional entropy loss function.
+We denote the $\epsilon$ obtained from minimizing (18) as $\epsilon_{CE}$.
+When $v_{k}(\epsilon)$ is unavailable in closed form, this loss function is estimated from the posterior samples using (9).
+For difference boundary analysis on a set $L$ containing $K$ pairs of neighboring regions and a corresponding set of difference probabilities, $\{v_{ij}(\epsilon)\}_{(i,j)\in L}$, the summation in (18) is over all pairs of neighboring regions, $(i,j)\in L$.
+In the special case of perfect spatial health equity where all difference probabilities are exactly equal for all $\epsilon$, this procedure results in $v_{(i,j)}(\epsilon_{CE})=0.5$ for all boundaries $(i,j)$.
+Also, $v_{k}(\epsilon)$ is replaced with $h_{k}(\epsilon;\rho)$ defined in (10) when conditioning on a given $0<\rho<1$.
+As an alternative to optimizing (18), we can consider the number of positive decisions indicated by $T$.
+Fixing two quantities out of $\epsilon$, $\delta$ and $T$ determines the third, representing a three-way trade-off between protection against Type I errors, quantity, and quality in our positive results: $\delta$ indicates our tolerance for false positives; $T$ is the quantity of positive results; and $\epsilon$ is our standard for a disparity, which reflects the quality of our decisions.
+If $\delta$ and $T$ are fixed, then a valid choice of $\epsilon$ should belong to $A_{\epsilon}=\left\{\epsilon>0:\overline{\text{FDR}}\left(t^{\star}(\epsilon)\right)\leq\delta,\;\sum_{(i,j)\in L}d_{ij}\geq T\right\}$.
+This criterion can be more useful in policy settings aiming to identify rankings of the most egregious disparities, for example, the top fifty.
+However, this approach will always report $T$ disparities, even if there are no significant differences between the regions.
+Finally, by fixing $\epsilon$ and $T$, $\delta$ implicitly equals the Bayesian FDR of the resulting decisions, although this does not represent active protection against Type I errors.
+We conclude this section with a remark on the implications of the limiting cases discussed in Section 2.2.
+As $\rho$ tends to 0, (6) indicates that boundary detection becomes meaningless because spatial effects converge to 0.
+As $\rho$ tends to 1, $\sigma^{2}$ and $\gamma$ also have limiting distributions described in (7), suggesting that the conditional variance of $\gamma$ (in the definition of the $\epsilon$-difference probability) or $\sigma^{2}$ will neither explode nor shrink toward 0 when $\rho$ is near 1.
+4. Simulation Experiments
+4.1. Simulation Comparison with ARDP-DAGAR
+We compare difference boundary classification performance between our $\epsilon$-difference approach and the areally-referenced Dirichlet process (ARDP) method introduced by Li et al.
+(2015) and expanded to the multivariate setting in Gao et al.
+(2023) in 100 simulated datasets on a map of the Californian counties.
+Figure 1.
+Left: simulated unscaled spatial effects $\phi$.
+There are 90 difference boundaries separating counties with distinct spatial effects.
+Right: prior density of $\rho\sim\mbox{PC}(\lambda_{\rho}=0.2)$.
+To create a reference dataset, we first draw $z\sim\mbox{N}_{n}(0_{n},\Sigma)$, where $\Sigma$ is a exponential covariance kernel such that $\Sigma_{ij}=\exp(-d_{ij}/150)$ where $d_{ij}$ is the Great Circle distance between the centroids of regions $i$ and $j$ in kilometers.
+We divide the components $z_{1},\ldots,z_{n}$ into quintiles and replace each component with the mean of the corresponding quintile.
+This neutral setup allows for a fair comparison between our proposed method and an existing Bayesian non-parametric approach.
+The resulting spatial effects $\phi$ are fixed across datasets and mapped in the left panel of Figure 1.
+Out of 139 total, there are 90 boundaries $(i,j)$ where $\phi_{i}\neq\phi_{j}$, which we label as the true difference boundaries.
+We then generate 100 datasets $\{X^{(i)},y^{(i)}\}$ by generating covariates $X^{(i)}=(1_{n},x^{(i)}),x^{(i)}\sim\mbox{N}_{n}(0,I_{n})$ and the response $y^{(i)}\sim\mbox{N}_{n}(X^{(i)}\beta+\gamma,\sigma^{2}(1-\rho)I_{n})$ with $\gamma=\sqrt{\sigma^{2}\rho}\phi$, $\rho=0.95$, $\sigma^{2}=5$, and $\beta=(2,5)^{\mathrm{\scriptscriptstyle T}}$.
+For both methods, we compute difference probabilities over $L$, the set of all boundaries between Californian counties.
+For the $\epsilon$-difference method, we employ the BYM2 model in (1) with the prior $\pi(\beta,\gamma,\sigma^{2},\rho)=\mbox{N}_{n}(\gamma\,|\,0_{n},\sigma^{2}\rho V_{\phi})\times\mbox{IG}(\sigma^{2}\,|\,0.001,0.001)\times\mbox{Unif}(\rho\,|\,0,0.99)$.
+We use a spatial CAR prior with $V_{\phi}^{-1}=c(D_{W}-\alpha W)$ constructed using the adjacency relations of the Californian counties, $\alpha=0.99$, and $c=0.8352$ such that the geometric mean of the marginal prior variances of $\gamma$ is equal to one.
+We choose $\lambda_{\rho}=0.2$, corresponding to a prior belief that $\mathbb{P}(\rho\leq 0.5)\approx 0.67$ and truncate the support of $\rho$ to the interval $[0,0.99]$ to improve convergence.
+The density of $\mbox{PC}(\rho\,|\,\lambda_{\rho}=0.2)$ is shown in Figure 1.
+For each dataset, we draw $20,000$ posterior samples of $(\beta,\gamma,\rho,\sigma^{2})$ after 40,000 burn-in samples using the Metropolis within Gibbs sampling algorithm detailed in Appendix LABEL:sec:bym2_sampling_algorithm, obtain $\epsilon_{CE}$ by minimizing (18), and estimate the difference probabilities $\{v_{ij}(\epsilon_{CE})\}_{(i,j)\in L}$ using (12).
+The ARDP model builds spatial dependence through a Markov random field on a latent component and imbues the spatial effects with a probability distribution.
+Following the notation of Li et al.
+(2015), we write the ARDP model as
+[…]
+where $F^{(i)}(\cdot)$ is the CDF of the marginal component $\phi_{i}$, $u_{1},\ldots,u_{n}$ are indices sampled from $\{1,\ldots,N\}$, and $\{p_{1},\ldots,p_{K}\}$ are the stick-breaking weights truncated to $K$ clusters.
+The spatial components $\phi$ follow a directed acyclic graphical autoregression (DAGAR) model parameterized by $r\in(0,1)$, a spatial autocorrelation parameter (Datta et al., 2019).
+The DAGAR model requires a specific ordering of the regions $\{1,\ldots,n\}$, which we construct by sorting $\delta_{1}<\delta_{2}<\ldots<\delta_{n}$ where $\delta_{i}=L_{i,1}-L_{i,2}$ and $(L_{i,1},L_{i,2})$ are the coordinates of a Albers projection of the $i$th county’s centroid.
+Denote $N(i)$ as the set of neighboring regions that precede $i$ in the ordering and $a_{i}$ is the number of regions in $N(i)$.
+Then, $W_{\phi}^{-1}=(I_{n}-B)^{\mathrm{\scriptscriptstyle T}}R_{\phi}(I_{n}-B)$, where $B_{ij}=0$ if $j\notin N(i)$, $B_{ij}=\frac{r}{1+(a_{i}-1)r^{2}}$ for $i=2,\ldots,n$ and $j\in N(i)$, and $R_{\phi}$ is a diagonal matrix with $(R_{\phi})_{ii}=\frac{1+(a_{i}-1)r^{2}}{1-r^{2}}$ for $i=1,\ldots,n$.
+We match the simulation settings in Gao et al.
+(2023) by setting the prior $\pi(\beta,\sigma^{2},\sigma^{2}_{s},r)=\mbox{N}_{p}(\beta\,|\,0_{p},100^{2}I_{p})\times\mbox{IG}(\sigma^{2}\,|\,2,0.1)\times\mbox{IG}(\sigma^{2}_{s}\,|\,2,1)\times\mbox{Unif}(r\,|\,0,0.99)$ and fixing $\alpha_{V}=1$.
+For each simulated data set, we draw 50,000 $\times$ 4 chains = 200,000 samples from the joint posterior $\pi(\beta,\gamma,\sigma^{2},\sigma^{2}_{s},\phi,r\,|\,y)$ after 50,000 burn-in samples per chain using the MCMC algorithm implemented in Gao et al.
+(2023).
+We use the posterior samples of $\gamma$ to compute Monte Carlo estimates of the difference probabilities $w_{ij}=\mathbb{P}(\gamma_{i}\neq\gamma_{j}\,|\,y)$ for all neighboring pairs $(i,j)\in L$.
+We also compare our results to spatial outlier detection with Local Moran’s I computed with “rook contiguity” weights based on the standardized residuals $z_{i}=e_{i}-\frac{1}{n}\sum_{j}e_{j}$ from an ordinary least squares regression of $y^{(i)}$ on $x^{(i)}$.
+We compute significance p-values for each region using 50,000 permutations and control the FDR at $0.15$ with the classical method (Benjamini and Hochberg, 1995).
+If a region is significant after multiplicity adjustment, it is a high-low outlier if $z_{i}>0,\;\sum_{j\,:\,i\sim j}z_{j}<0$ or a low-high outlier if $z_{i}<0,\;\sum_{j\,:\,i\sim j}z_{j}>0$.
+We report $(i,j)$ as a difference boundary if the region $i$ is a high-low or low-high outlier and the region $j$ is not significant or is not of the same type of outlier.
+Figure 2.
+Classification performance of $\epsilon$-difference method with CAR spatial kernel and ARDP-DAGAR method for data simulated with $\rho=0.95$ (top panels) and $\rho=0.7$ (bottom panels).
+Left: ROC curves showing true positive rate versus false positive rate colored by method across varying number of reported difference boundaries.
+Performance at each cutoff value is averaged across 100 simulated datasets.
+Middle: violin plots with of difference probabilities generated across each dataset and method.
+Difference probabilities corresponding to true simulated disparities are colored in blue.
+Right: histograms of the number of reported disparities in each dataset colored by method.
+The Bayesian FDR tolerance is set to $\delta=0.3$ for the ARDP-DAGAR and $\epsilon$-difference methods while the FDR-adjusted p-value significance threshold is set to $0.15$ for boundary detection using Local Moran’s I.
+The difference probabilities $v_{ij}(\epsilon)$ and $w_{ij}(\epsilon)$ have equivalent interpretation: a higher difference probability indicates a stronger posterior belief that $(i,j)$ is a true difference boundary.
+For $T=1,\ldots,139$, we take the boundaries corresponding to the top $T$ difference probabilities as the reported difference boundaries in each dataset and average sensitivity and specificity across datasets for each method to obtain the averaged receiver operating characteristic (ROC) curve in the top left panel of Figure 2.
+For all values of $T$, our proposed framework shows comparable performance to the nonparametric alternative.
+The area under the ROC curve (AUC) for the $\epsilon$-difference method is 0.856, while the AUC for the ARDP-DAGAR method is 0.871.
+When $T=90$, the $\epsilon$-difference method has an averaged sensitivity of 0.828 and a specificity of 0.683, while the ARDP-DAGAR method has an averaged sensitivity of 0.832 and specificity of 0.691.
+We apply the FDR control in Section 3.4 with tolerance $\delta=0.3$ for each data set and both types of difference probabilities.
+The top difference probabilities from ARDP-DAGAR in the top middle panel of Figure 2 belong to a few datasets, resulting in a large variation in the number of reported disparities, as shown in the right panel of Figure 2.
+The ARDP-DAGAR produced 58.59 (SD = 29.71) disparities reported on average, an overall false discovery rate of 10.68%, and 6 out of 50 datasets with zero reported disparities.
+Boundary detection based on Local Moran’s I reported 11.87 (SD = 6.63) disparities per data set on average, an overall false discovery rate of 4.30%, and 10 datasets reporting no disparities.
+The $\epsilon$-difference approach reported 19.54 (SD = 8.45) disparities on average, an overall false discovery rate of 0.92%, and zero data sets with no reported disparities.
+Therefore, our approach shows an attractive decrease in the probability of reporting global non-significance while minimizing the false discovery rate.
+The performance of all three methods suffer greatly when the spatial signal-to-noise ratio decreases.
+We repeat the simulation runs with $\rho=0.7$, and re-evaluate the aforementioned performance metrics in the bottom panels of Figure 2.
+The AUC for the $\epsilon$-difference framework decreases to 0.684 and the AUC for the ARDP-DAGAR method decreases to 0.669.
+In both cases, our proposed framework on average achieves comparable classification performance to the ARDP-DAGAR approach and results in fewer datasets with zero reported disparities under identical constraints on the Bayesian FDR.
+Boundary detection using Local Moran’s I resulted in a false discovery rate of 18.67%, while the $\epsilon$-difference (8.97%) and ARDP-DAGAR (18.42%) methods provided greater control over the FDR.
+Because of the discrete clustering property of Dirichlet process models, our approach performs better when the spatial residual surface is smooth and spans a wide range of values, while the ARDP-DAGAR method is more suitable when the spatial residuals fall into a small number of distinct levels.
+Our method may be more efficient for approximately normally distributed outcomes, in particular disease maps in which only one aggregated incidence or prevalence rate is recorded for each areal unit.
+Compared to spatial outlier detection methods such as those based on Local Indicators of Spatial Association, our approach offers full Bayesian model estimation and FDR control in the presence of non-spatial heterogeneity.
+4.2. Simulated Count Outcome
+We demonstrate our analysis framework for a simulated count outcome on a map of the $n=58$ Californian counties.
+We simulate a latent risk factor $\gamma\sim\mbox{N}_{n}(0_{n},\sigma^{2}\rho W_{\phi})$ and unstructured error $\eta\sim\mbox{N}_{n}(0_{n},\sigma^{2}(1-\rho)I_{n})$ with $\sigma^{2}=2,\rho=0.93$, and $W_{\phi}$ as an exponential covariance kernel such that $W_{\phi,ij}=\exp(-d_{ij}/50)$, where $d_{ij}$ is the great circle distance between the centroids of county $i$ and $j$ in kilometers.
+We generate covariates $x_{i}=(1,x_{i1})^{\mathrm{\scriptscriptstyle T}},x_{i1}\sim\mbox{N}(2,1)$ and exposure $E_{i}=\lceil H_{i}\rceil,H_{i}\sim\text{Unif}(10^{4},5\cdot 10^{4})$ for $i=1,\ldots,58$.
+We simulate $y_{i}\sim\text{Poisson}(E_{i}\exp(x_{i}^{\mathrm{\scriptscriptstyle T}}\beta+\gamma_{i}+\eta_{i}))$ with $\beta=(-5,0.5)^{\mathrm{\scriptscriptstyle T}}$.
+Here, we define the simulated rate as $y_{i}/E_{i}$ and denote the boundary between county $i$ and $j$ as a true spatial disparity at the $\epsilon$-difference level if $\frac{|\phi_{i}-\phi_{j}|}{a_{ij}^{\mathrm{\scriptscriptstyle T}}W_{\phi}a_{ij}}>\epsilon$.
+We analyze the simulated data using a Poisson likelihood with the link function in (13) as $g(y_{i})=\log\left(\frac{y_{i}}{E_{i}}\right)$ and use the prior $\pi(\beta,\gamma,\eta,\sigma^{2},\rho)\propto\text{N}_{n}(\gamma\,|\,0_{n},\sigma^{2}\rho V_{\phi})\times\text{N}_{n}(\eta\,|\,0_{n},\sigma^{2}(1-\rho)I_{n})\times\text{IG}(\sigma^{2}\,|\,0.1,0.1)\times\mbox{PC}(\rho\,|\,0.2)$.
+We place a CAR spatial prior on $\gamma$ by fixing $V_{\phi}^{-1}=c(D_{W}-\alpha W)$ using the adjacency relations of the Californian counties, $\alpha=0.99$, and $c=0.8352$.
+Using Hamiltonian Monte Carlo implemented within the rstan package in R, we obtain $20,000\times 4\;\text{chains}=80,000$ total posterior samples of $(\beta,\gamma,\sigma^{2},\rho)$ after discarding 40,000 initial burn-in samples per chain.
+We consider $L$, the set of boundaries between two neighboring California counties, and use Monte Carlo estimates of $\{\tau_{ij}(\epsilon)\}_{(i,j)\in L}$ via (14) to minimize the conditional entropy loss and obtain the optimal difference threshold $\epsilon_{CE}=0.738$.
+Of 139 total boundaries, 76 are a true disparity at the $\epsilon_{CE}=0.738$ difference level.
+We repeat the analysis with two alternative difference probability definitions: $\psi_{ij}=\mathbb{P}(|\phi_{i}-\phi_{j}|>\epsilon\,|\,y)$, based on the unstandardized log-rate difference, and $\theta_{ij}=\mathbb{P}(\exp(|\gamma_{i}-\gamma_{j}|)>\epsilon\,|\,y)$, based on the adjusted incidence rate ratio (IRR).
+For a fair comparison, we also define a pair $(i,j)$ to be a true $\psi$-disparity on the $\epsilon$-difference level if $|\phi_{i}-\phi_{j}|>\epsilon$ or a $\theta$-disparity if $\exp(|\gamma_{i}-\gamma_{j}|)>\epsilon$.
+We minimize the conditional entropy loss function to obtain $\epsilon_{1}=0.465$ and $\epsilon_{2}=1.506$ for the $\psi_{ij}$ and $\theta_{ij}$ difference thresholds, respectively.
+Figure 3.
+Boundary detection results for simulated count outcome on a California county map.
+From left to right: (1) Map of simulated rates $y_{i}/E_{i}$.
+(2) Map of simulated latent factors $\gamma$ with (1) boundaries associated with top 20 $\epsilon_{CE}=0.738$ standardized log-rate difference probabilities $\tau_{ij}$ and (3) boundaries associated with top 20 $\epsilon=1.506$ IRR difference probabilities $\theta_{ij}=\mathbb{P}(\exp(|\phi_{i}-\phi_{j}|)>\epsilon\,|\,y)$.
+(4) ROC curves showing true positive rate versus true negative rate across varying cutoff values using $\tau_{ij}$ difference probabilities (red; AUC = 0.847), $\psi_{ij}=\mathbb{P}(|\phi_{i}-\phi_{j}|>\epsilon\,|\,y)$ difference probabilities (green; AUC = 0.733), and $\theta_{ij}=\mathbb{P}(\exp(|\gamma_{i}-\gamma_{j}|)>\epsilon\,|\,y)$ difference probabilities (blue; AUC = 0.723).
+Figure 3 shows maps of the simulated rates (left panel) and spatial residuals $\gamma$ with the boundaries corresponding to the top 20 log-rate difference probabilities $\tau_{ij}$ (middle left) or IRR difference probabilities $\theta_{ij}$ (middle right).
+The two types of difference probabilities share 13 out of the top 20 difference boundaries and both types resulted in all pairwise difference probabilities being below $0.70$.
+We evaluate the overall classification performance of the difference probabilities $\{v_{ij}(\epsilon)\}_{(i,j)\in L}$ in detecting spatial disparities and obtain a favorable classification performance shown by the ROC curve (AUC = 0.847) in the rightmost panel of Figure 3.
+We obtain an AUC value of $0.733$ when using $\psi_{ij}$ difference probabilities to detect $\psi$-disparities and an AUC value of $0.723$ when using $\theta_{ij}$ difference probabilities to detect $\theta$-disparities.
+To evaluate rank stability, we compute the Spearman correlation between $\{\tau_{ij}(\frac{1}{3}\epsilon_{CE})\}_{(i,j)\in L_{S}}$ and $\{\tau_{ij}(3\epsilon_{CE})\}_{(i,j)\in L_{S}}$ as $r_{s}=0.903$, which we consider as a rank stability score.
+The rank stability score using the unstandardized difference in $\psi_{ij}$ is 0.825, indicating that the proposed difference probability formulation $\tau_{ij}$ in (14) offers improved classification performance and robustness in the rankings of the top difference probabilities under this simulation example.
+5. Application
+We analyze publicly available health data from the Institute of Health Metrics and Evaluation (IHME) on age-standardized county-level estimates of mortality rates in tracheal, bronchus, and lung cancer in 2014 (Mokdad et al., 2017).
+We adjust for multiple risk factors by incorporating county-level estimates of smoking prevalence in both sexes in 2012, physical inactivity in 2014, percentage of residents 18 years or older without health insurance from 2012 to 2016, unemployment rate percentile (0-1.0) in 2014, overall social vulnerability index (SVI) percentile in 2014, adult diabetes prevalence in 2014, and adult obesity prevalence in 2014.
+Smoking prevalence was originally derived using data from the Behavioral Risk Factor Surveillance System in Dwyer-Lindgren et al.
+(2014), while data for all other risk factors was downloaded from the publicly available US Diabetes Surveillance System (Centers for Disease Control and Prevention, ).
+Previous work in Shreves et al.
+(2023) indicates that lung cancer mortality rates exhibit spatial autocorrelation that is not completely accounted for by smoking prevalence.
+For our study region, we subset into $n=3,017$ contiguous US counties with available estimates for smoking prevalence and lung cancer mortality rates with the aim of detecting disparities corresponding to county boundaries.
+Although we drop counties with missing data for the sake of simplicity in this application, one can refer to Rubin (1976) for an introduction to Bayesian approaches to missing data or Ma and Chen (2018) for a modern overview.
+For simplicity, we drop these counties from our study region as the remaining counties form a contiguous region.
+The cancer mortality rate and estimates of total smoking prevalence for the final study region are shown in Figure 4.
+High-risk regions can be clearly identified, most notably in the East South Central states, and our analysis aims to identify local health inequity at the county level.
+Figure 4.
+Left: map of county-level age-standardized mortality rate estimates for lung cancer in 2014, colored by quintile.
+Right: map of county-level total smoking prevalence estimates in 2012, colored by quintile.
+To assess spatial autocorrelation, we fit an ordinary least squares linear model with the aforementioned risk factors as predictors of lung cancer mortality before computing Moran’s $I=0.397$ and Geary’s $C=0.586$ using the residuals (Banerjee et al., 2015).
+Both statistics were significant with 10,000 random permutations (p-value = 1/10,001), suggesting that spatial autocorrelation is not fully captured by linearly regressing mortality rates on smoking prevalence.
+We employ (1) such that $y\,|\,\beta,\gamma,\sigma^{2},\rho\sim\text{N}_{n}\left(X\beta+\gamma,\sigma^{2}(1-\rho)I_{n}\right)$ where $\beta$ is the vector of regression coefficients with an intercept.
+Similarly to the simulation examples, we place a CAR prior on $\phi$ by setting $V_{\phi}^{-1}=c(D_{W}-\alpha W)$, where the neighbor and adjacency matrices are computed using the subset of US counties, $\alpha=0.99$, and $c=0.3762$.
+We treat $\rho$ as an unknown parameter and specify the joint prior as $\pi(\beta,\gamma,\sigma^{2},\rho)=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\text{PC}(\rho\,|\,\lambda_{\rho})$, where $a_{\sigma^{2}}=0.1$ and $b_{\sigma^{2}}=0.1$.
+Although the set of regions here differs slightly from the simulation example, we also set $\lambda_{\rho}=0.0335$ as it approximately satisfies $\mathbb{P}(\rho\leq 0.5)=\frac{2}{3}$.
+After 10,000 burn-in samples, we draw 30,000 samples from the posterior of $\{\beta,\gamma,\sigma^{2},\rho\}$ using Algorithm LABEL:alg:gibbs_sampling.
+Table 1 presents 95% credible intervals for the nonspatial effect parameters, indicating that smoking, SVI percentile, physical inactivity, and uninsured rate are significant predictors of lung cancer mortality.
+The credible interval of $\rho$ suggests that we are able to learn effectively about $\rho$, and the CAR spatial model explains most of the variance in the data.
+Parameter Description Posterior Mean 95% Credible Interval
+$\beta_{0}$ Intercept 9.56 (4.828, 14.223)
+$\beta_{1}$ Smoking prevalence (%) 1.90 (1.754, 2.052)
+$\beta_{2}$ Unemployment percentile 1.85 (-0.473, 4.174)
+$\beta_{3}$ SVI percentile 5.70 (2.755, 8.68)
+$\beta_{4}$ Physically inactive (%) 0.35 (0.202, 0.49)
+$\beta_{5}$ Uninsured (%) -0.60 (-0.746, -0.451)
+$\beta_{6}$ Diabetes prevalence (%) 0.18 (-0.137, 0.506)
+$\beta_{7}$ Obesity prevalence (%) 0.06 (-0.08, 0.2)
+$\sigma^{2}$ Total variance 108.25 (95.927, 123.349)
+$\rho$ Spatial proportion of variance 0.78 (0.692, 0.861)
+Table 1. Posterior summaries of non-spatial effect parameters in analysis of US county-level lung cancer mortality rates.
+We detect spatial disparities using the $\epsilon$-difference framework in Section 3.
+For this dataset, we denote the set of all $K=$ 8,793 pairs of neighboring counties as $L=\{(i,j):i<j,i\sim j\}$ and $\epsilon$-difference probabilities of the form $v_{ij}(\epsilon)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y\right)$ for all $(i,j)\in L$.
+We select the difference threshold $\epsilon$ by minimizing the conditional entropy loss in (18), which yields $\epsilon_{CE}=0.922$.
+For this analysis, we set a maximum allowable Bayesian FDR of $\delta=0.05$, compute $t^{\star}=0.884$ using (17) and report 722 county boundaries exhibiting spatial disparities.
+Figure 5.
+Map of posterior predictive means of lung cancer mortality rates.
+Red lines correspond to disparities between neighboring counties with difference threshold $\epsilon_{CE}=0.922$ and Bayesian FDR tolerance $\delta=0.05$.
+The posterior predictive means and detected disparities are mapped in Figure 5.
+The reported disparities are mainly in relatively high risk states (KY, GA, TN, WV, MS, MS, MO, VA), where several county mortality rate estimates are in the upper quintile with a wide range, suggesting prevalence of health inequities on the local county scale.
+These disparities between neighboring counties may be caused by a combination of local differences between counties’ health policies or other latent determinants of health.
+Although our model assumes a global effect from each risk factor, previous analysis by Dong et al.
+(2022) suggests that the association of risk factors such as obesity/diabetes prevalence and physical inactivity with cancer mortality may vary by geographical region.
+The complete list of 722 detected difference boundaries ranked by their corresponding difference probabilities is available in Table LABEL:tab:RDA_disparities_table.
+6. Discussion
+Our $\epsilon$-based difference boundary detection is considerably simpler to implement and interpret than Bayesian nonparametric approaches (Li et al., 2015; Gao et al., 2023).
+We are able to stabilize the difference probabilities with respect to $\epsilon$, which allows public health officials to report, say, the top 50 disparities between neighboring counties in a robust way.
+We could also define a difference boundary based on the posterior predictive distribution or the posterior distribution of the total residuals $(\gamma+\eta)$ rather than the spatial random effects to identify transitions in the original health outcome or deviations from the explanatory model.
+We provide some insights in Section LABEL:sec:model_response_diffs and establish stabilized rankings for the exact conjugate model.
+This may be preferred in public health practice because of its interpretation in terms of the health outcome itself, but such differences will conflate disparities attributed to space with nonspatial explanatory variables or risk factors.
+Future areas of investigations will include the impact of spatial confounding on $\epsilon$-based boundary detection, and extending our framework to multivariate modeling with dependent outcomes.
+For computational efficiency, we will also develop Bayesian predictive stacking (as developed in Zhang et al., in press; Pan et al., 2025, for Gaussian process models) to avoid MCMC iterations and achieve exact Bayesian boundary detection by stacking over conjugate Bayesian models corresponding to different values of $\rho$.
+Supplementary Materials
+Supplementary materials contain derivation of rank-stability results, sampling algorithm details, formulation of an $\epsilon$-difference probability on the modeled response level, simulation analyses of classification performance sensitivity to the choice of $\epsilon$ and uncertainty in $\rho$, and the full list of spatial disparities in the data analysis.
+Computer programs that reproduce results in this manuscript are available at https://github.com/Ky-Wu/bayesian_spatial_health_disparities.
+Algorithms LABEL:alg:gibbs_sampling and LABEL:alg:update_steps are executed in R using RcppArmadillo, the ARDP-DAGAR approach is executed using rjags, and MCMC sampling for the BYM2 Poisson model in Section 4.2 is performed using rstan.
+This work used computational and storage services associated with the Hoffman2 Shared Cluster provided by the UCLA Office of Advanced Research Computing’s Research Technology Group.